# Robust Capacity Expansion Modelling for Renewable Energy Systems under Weather Uncertainty

### v3

@@
Robust Capacity Expansion Modelling for Renewable Energy Systems under Weather Uncertainty
Sebastian Kebrich Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany Correspondence: s.kebrich@fz-juelich.de These authors contributed equally Felix Engelhardt RWTH Aachen University, Faculty of Computer Science, Teaching and Research Area Combinatorial Optimization, Aachen, Germany Correspondence: engelhardt@combi.rwth-aachen.de These authors contributed equally David Franzmann Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany Christina Büsing RWTH Aachen University, Faculty of Computer Science, Teaching and Research Area Combinatorial Optimization, Aachen, Germany Jochen Linßen Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany Heidi Heinrichs Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany University of Siegen, Department of Mechanical Engineering, Chair for Energy Systems Analysis, Siegen, Germany
@@ INTRODUCTION
Modelling energy systems requires access to different types of data. For future greenhouse gas neutral energy systems, weather time–series data is of special importance, as a vast expansion of wind power and solar photovoltaic (PV) that are dependent on the weather is perceived as indispensable 1, 2. However, we can not expect to have accurate long–term information on future weather conditions. Instead, historical time–series serve as a substitute in capacity expansion planning, and often a single ”representative” sample year is selected for optimisation 3. For example, the International Renewable Energy Agency (IRENA) recommends using $2018$ as a reference year because it represents generation from renewable technologies well on average 4. In the following, we argue that using ”representative” sample years may lead to energy system models (ESMs) that appear sound at first glance, but would fail to meet supply in reality. We then propose a tractable approach to counteract this effect and to make energy systems more robust against (weather) uncertainty, i.e. ensure security of supply.
@@
There is ample evidence in literature that the results of energy system optimisation are sensitive to changes in weather time–series data. Schyska et al. (2021) evaluate the sensitivity of capacity expansion models with regards to multiple sources of uncertainty. They conclude that some years are unsuited as reference years, as using them for optimisation leads to significant misallocation of assets.5 The works by Haddeland et al. (2022) 6, and Staffel et al. (2018)  7, which builds on earlier work by Pfenninger et al. (2017) 8, similarly find the choice of weather years significantly effects renewable generation and thus power output. Not only that, but the effect of the weather increases with increasing share of renewable technologies 9. Not all research agrees on this; Schlachtenberger et al.(2018) 10 optimise three weather years with hourly data both individually and as one time–series with a resolution of 3h per time step, finding only small variations in total annual cost (TAC) and installed capacities. However, they note that aggregating multiple hours together introduces a smoothing effect that systematically favours photovoltaic and underestimate battery and wind generation requirements.
@@
Additionally, there is a fundamental tension between identifying typical years, and build energy system models that are protected against extreme events, i.e. robust. On this, Hilbers et al. (2020) 11 introduce a method of importance subsampling for time–series aggregation to explicitly preserve extreme events in the weather data as an alternative to established ”representative days” clustering approaches. Ryberg’s dissertation 12 investigates the impact of generation lulls in a energy system for a large part of Europe calculating backup capacities required to overcome these. Ruhnau et al. (2022) 13 look into the storage requirements for a renewable–based energy system model for Germany using $35$ years of weather data taking consecutive extreme events into account. They conclude that consecutive extreme events increase storage requirements significantly compared to even the most extreme, but singular events. Thus, atypical weather time–series may be particularity well-suited for optimisation because they capture important system behaviours, e.g. dark lulls, that significantly impact energy system models. In this context, Grochowicz et al. (2024) 14 discuss optimising sequential weather years. They use a geometry–based solution approach targeting the solution space. In follow–up work, they use electricity shadow prices to identify difficult weather periods 15. They observe that such difficult weather periods are not just meteorological events, but results of the interplay of meteorology and electricity storage and network structures.
@@
In stochastic optimisation, the notion of using typical, i.e expected, behaviour as a baseline for optimisation is a well known concept. Its usefulness is determined by the value of a stochastic solution, which represents the gap in expected performance of a solution obtained with expected data and one determined by solving a full stochastic optimisation problem  16. In general, this gap may be arbitrarily large. However, research has also shown that we can still extract insights from solutions obtained with with such data  17. Specifically concerning weather robustness, Forghani et al. (2025) propose an intermediate approach where not only a representative but also worst-/best-cost years are used as input data for a stochastic optimisation. Their results show that this reduces loss-of-load to practically zero, at $1\%$ additional cost 18.
An important modelling decision in this context is the value and availability of recourse, i.e. the ability to specify or change parts of the solution if uncertainty realises. In capacity expansion planning, imports constitute such recourse. If arbitrary imports are allowed, any misplanning can be compensated for, if at a cost. We focus on a different setting, the one where imports are limited or energy systems are to be self–sufficient. Here, we use adaptive robust optimisation (ARO) to model limited recourse while still enforcing strict guarantees on security of supply.
@@
ARO, alternatively called two-stage robustness, is characterised though a bilevel structure 19 in which a decision maker has to make a set of first stage decisions $x\in\mathcal{X}$, then one of multiple scenarios $u\in\mathcal{U}$ may realise and afterwards the decision maker has the option to react to scenario $u\in\mathcal{U}$ with a scenario-dependent set of decisions $y(u)\in\mathcal{Y}(x,u)$. Equation (ARO) illustrates the structure of a ARO problem. Here, $c$ and $d$ are vectors of first and second stage costs, respectively. The matrix $A$ encodes all information that is not affected by uncertainty, whereas the matrices $B(u)$ and the right-hand side vector $b(u)$ are dependent on the scenario $u$.
@@
In the setting of this work, the first stage decisions correspond to the capacity expansion planning, e.g. investment in solar power plants, power lines, storage units, and the second stage corresponds to operational decisions, e.g. how much energy is to be stored in which battery at what time. If energy imports are allowed, these will also be encoded in $\mathcal{Y}$. The scenario $u$ corresponds to a specific weather year. Table 1 illustrates this setting. Here, we consider different locations $A,B,C$ for a solar power plant, whose production is subject to changing weather patterns in weather years $u_{1},u_{2}$.
@@
Table 1: Investment problem: Chose number of solar power plants to build on locations $A,B,C$. Values indicate expected energy supply in GWh/year per investment.
@@
The optimal solution depends on what type of recourse is available and how much supply is needed. For example, if we assume similar cost at all locations and no/expensive imports, investing in location $B$ is the most cost effective for meeting a demand of 7575MWh/year. However, if the total demand were 200200MWh/year, it is more effective to invest in $A$ and $C$ instead of $B$.
Next, we look at general solution approaches. To the best of our knowledge, Zeyringer et al. (2018) 20 were the first to assess the effects of weather uncertainty in input data on capacity–expansion planning, looking at an energy system model of Great Britain. They propose using multiple historical time series, optimising over each, and then evaluating the operational costs/supply costs incurred by the proposed energy systems under the assumption that the time–series of a different historical year realises. Then, an energy system model with lowest worst–case costs/lowest supply gap amongst the solutions is selected. They find that starting with the ‘’wrong” reference year may lead to misallocation of resources and that modelling with multiple years leads to more consistent results with lower worst–case cost, the latter of which had also been noted before 21. For the US, Dowling et al. (2020) 22 make the case for multi–year modelling to accurately capture long–term storage effects, noting that the cost of variable renewable power systems are especially sensitive to long–duration storage costs. A recent work by Gøtske et al. 23 also assesses energy systems based on different weather years. They employ CO2 emitting backup technologies, and analyse structural elements of the respective solutions. Either approach allows for selecting potentially more suitable energy system models, but it can not assure solutions meet certain supply/demand across all years.
@@
Returning to our example in Table 1, the approach by Zeyringer et al. would not be able to find the optimal robust solution for a total demand of 7575MWh/year, which is to invest in location $B$. This is because if the future is known, i.e. we assume either $u_{1}$ or $u_{2}$ to realise, the optimal investment is always to either invest in location $A$ or location $B$. An alternative approach was put forward by Caglayan et al. (2019) 24, who noted variations of up to $20\%$ variations in TAC, and significant differences in energy system designs when optimising an integrated electrical/hydrogen ESM for Western Europe with different weather years. They propose an iterative approach, where they begin with optimising individual years, but then average their designs and take the average as a lower bound for installed capacities in a new round of optimisation 25. Since the (average) capacities installed during sequential algorithm iterations will be non-decreasing and bounded, convergence and thus a feasible solution will be guaranteed in most cases. In our example, the algorithm would enforce capacity >0>0 for location A and C, which would eventually lead to a feasible but unnecessarily expensive solution that uses both A and C.
@@
The algorithm proposed in this work consists of four main steps, as shown in Figure 1. As its input, the algorithm takes a set of $n$ weather times series corresponding to weather years, one of which is a designated starting/reference year. We begin with solving the capacity expansion planning problem for this year (CAPEX), then we evaluate the performance of the proposed investment in each of the other weather years by solving the corresponding unit commitment problem (UC), which formalises the validation proposed in previous works  20, 23.
To ensure feasibility, existing approaches allow for energy imports. These directly correspond to supply gaps, which we denote by $\delta$. If the supply gaps are sufficiently small, we terminate. Otherwise, instead of using supply gaps $\delta$ merely as a measure for solution quality, we propose using the information encoded in $\delta$ to iteratively make solutions more robust. For that, the optimisation problem is modified by either MOD 1 / Demand increase, MOD 2 / Synthetic time–series or MOD3 / Combine, before reoptimising CAPEX evaluating the loss of load again. We denote the modified base problem by CAPEX.
@@
Start: Solve CAPEX for reference year Solve UC for other n1n-1 years Small loss of load? Apply MOD1-3 to CAPEX Reoptimise modified CAPEX End: Robust energy system design yesno
@@
The main contributions of our work include not only insights derived for the specific energy system model, and an algorithm for solving CAPEX problems to find solutions that are robust to weather variations, but foremost a systematic comparison of different novel and existing approaches for designing robust energy system models. Based on this, we recommend three working modifications that all lead to robust energy systems systems.
@@
All approaches generate robust solutions, i.e. solutions with no supply gaps, subject to numerical tolerances. The performance of the different modifications is summarised in Table 2. The results show a moderate cost increase for robust models of $2-2.9\%$ compared to the highest total annual cost of a single year, which is a lower (dual) bound on the cost of a robust solution and costs of 113.6113.6bn€ annually for a detailed energy system model implemented in ETHOS.FINE that is to be as realistic as possible and 59.759.7bn€ annually for a simplified gurobipy model used for comparing the different approaches.
Modification Increased Invest Average Least
MOD 1 / Demand increase Li-Ion, CCGT, Salt caverns +5.7% +2.0%
Smoothed MOD 1 / Demand increase CCGT, Salt caverns +9.2% +2.4%
MOD 2 / Synthetic time–series gurobipy PV, CCGT, Salt caverns, Electrolysers +2.5% +1.6%
MOD 2 / Synthetic time–series ETHOS.FINE PV, CCGT, Salt caverns, Electrolysers +3.7% +2.9%
MOD3 / Combine PV, Wind, CCGT, Salt caverns, Electrolysers +10.6% +2.5%
Table 2: Comparison of the different modifications in terms of convergence and performance. Costs in bn€ per year (TAC). Average and least cost increase compared to best lower (dual) bound, the highest total annual cost of a single unmodified year.
Results for Optimising Individual Years
The total annual cost for energy systems for the $38$ node model of Germany within ETHOS.FINE for the $40$ different weather year time–series deviates around an average of 106106bn€ with between 96.496.4bn€  and 113.6113.6bn€ annually, which equals $9\%$ to $7\%$ compared to the default reference year $2018$ recommended by IRENA 4. The results of optimising each year independently are given by Figure 2. While the variations in overall TAC are limited, the energy system designs show substantial deviations. The cost shares of single technologies across the $40$ different single weather years vary by $69\%$ for hydrogen pipelines, $57\%$ for hydrogen salt caverns, $53\%$ for CCGT, $44\%$ for Li–ion batteries, $40\%$ for PV, $38\%$ for electrolysers, $22\%$ for the electricity grid, and $20\%$ for onshore wind, making it challenging to draw recommendations for planned capacity expansion for future energy systems.
Figure 2: Total annual cost comparison by technology for energy system models optimised from 1980–2019 aggregated for a $38$ node Germany model set up in ETHOS.FINE.
@@
Figure 3: Total annual cost comparison by technology for energy system models optimised from 198020191980-2019 on a single node Germany model set up using gurobipy.
@@
In summary, even small input data differences may lead to large variations in investment into different technologies. For example, offshore wind is not utilised at all in some years, but makes up over $13\%$ of the total annual cost in $2014$.
Note that the single node model is not directly comparable to the ETHOS.FINE model, as it contains various simplifications, i.e. it is a linear relaxation that does not contain any transmission cost. As such, solutions are significantly cheaper, with an average of 55.455.4bn€ with between $52.1$ and 59.759.7bn€ annually, i.e. $7\%$ to $6\%$ compared to the average.
Evaluating the Cost of Robust Solutions
The results for single years imply upper (primal) and lower (dual) bounds for the cost of robust solutions. The dual bound is given by the maximum total cost for a single year, since the cost of a robust solution cannot be lower than the cost of a single year. A primal bound is given by the maximum capacity for each technology per region and year, since this will be feasible for all weather years. This results in a possible range of [113.6,195.66][113.6,195.66]bn€ annually for an optimal robust energy system in ETHOS.FINE. For the gurobipy model, the range for optimal robust solutions is [59.7,67.6][59.7,67.6]bn€ annually.
As shown in Figure  2, the proposed robust solutions are close to the dual bound, i.e. require minimal additional costs.
Results for Feasibility Testing
Feasibility testing shows that all $40$ energy system designs solely based on one year’s time–series lead to supply gaps in multiple other years. This means that none of the system designs are robust. This strongly supports the results by Cagalayan et al. 24, 25 – weather patterns matter and ignoring in energy systems with significant renewables may lead to systems that have large supply gaps under anything but the most optimal conditions.
@@
Of the $40$ energy system designs the five most expensive years $1987,1996,2009,2010$ and $2013$ were selected to make them robust using MOD 2 / Synthetic time–series in the ETHOS.FINE model. These are hereon referred to as the five reference years. Figure 12 in the appendix shows where in the five selected reference years supply gaps would occur. Figure 4 illustrates the reasons for that by showing three uncritical (top row) periods with sufficient supply as well as three critical (bottom row) time periods with significant supply gaps. The periods have varying duration and they are identified via clustering and feasibility testing. Uncritical time periods are characterised by high availability of PV or onshore wind or both, while critical ones are characterised by low availability of PV and low to negligible onshore wind. Offshore wind plays only a minor role due to its limited utilisation. The most critical time period takes place in $1994$, see Figure 4f). Here, wind and solar supply indicate a dark lull. Combined with the cold period identified in the electricity data, this constitutes a cold dark lull.
@@
Figure 4: 66 time periods from the $40$ years of weather and 11 year of future electricity demand data for Germany. The electricity demand is normalised to prevent overweighing; the weather data is aggregated. The upper three diagrams represent non–critical time periods, the lower three critical ones. Subfigure a) is a typical summer period with high PV availability and low electricity demand due to low heating requirements. In b) and c), typical autumn and winter period are shown. They are characterised by low availability of PV, but ample wind power to supply sufficient electricity. Note the increased electricity demand due to increased heating required. In d), a typical dark lull is characterised by low availability of PV and negligible amounts of wind, which coincides with high electricity demand due to increased heating. Subfigure e) shows an elongated dark lull period. Low availability of both PV and wind combined with increased electricity demand lead to overall difficult period requiring large amounts of hydrogen to be burned in the energy system. The last graphic f) shows the most critical period in the 40 years of weather data. Negligible wind combined with low availability of PV coincide with the highest electricity demand in the data due to high heating demand during an extreme cold spell in all of Germany.
Effect of Modifications
@@
All modifications lead to increased investment in CCGT, salt caverns and electrolysers, although the latter is less pronounced for MOD 1 / Demand increase, as seen in Table 2. This is expected, as CCGT provides weather independent energy supply and installing more CCGT also requires more H2 infrastructure such as salt caverns for storage and electrolysers for H2 conversion.

### v4

# arXiv:2504.06750v4[math.OC] 27 Feb 2026

@@
arXiv:2504.06750v4[math.OC] 27 Feb 2026
Robust Capacity Expansion Modelling for Renewable Energy Systems
Sebastian Kebrich Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany Correspondence: s.kebrich@fz-juelich.de, lead contact These authors contributed equally Felix Engelhardt RWTH Aachen University, Faculty of Computer Science, Teaching and Research Area Combinatorial Optimization, Aachen, Germany Correspondence: engelhardt@combi.rwth-aachen.de These authors contributed equally David Franzmann Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany Christina Büsing RWTH Aachen University, Faculty of Computer Science, Teaching and Research Area Combinatorial Optimization, Aachen, Germany Jochen Linßen Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany Heidi Heinrichs Forschungszentrum Jülich GmbH, Institute of Climate and Energy Research, Jülich Systems Analysis (ICE–2), Jülich, Germany University of Siegen, Department of Mechanical Engineering, Chair for Energy Systems Analysis, Siegen, Germany
@@
Future greenhouse gas neutral energy systems will be dominated by renewable energy technologies providing variable supply subject to uncertain weather conditions. For this setting, we propose an algorithm for capacity expansion planning: We evaluate solutions optimised on a single years’ data under different input weather years, and iteratively modify solutions whenever supply gaps are detected. These modifications lead to solutions with sufficient capacities to overcome periods of cold dark lulls and seasonal demand/supply fluctuations. A computational study on a German energy system model for $40$ operating years shows that preventing supply gaps, i.e. finding a robust system, increases the total annual cost by $1.6-2.9\%$. In comparison, non-robust systems display loss of load close to $50\%$ of total demand during some periods. Results underline the importance of assessing the feasibility of energy system models using atypical time-series, combining dark lull and cold period effects.
@@ INTRODUCTION
Modelling energy systems requires access to different types of data. For future greenhouse gas neutral energy systems, time–series data is of special importance, as a vast expansion of wind power and solar photovoltaic (PV) that are dependent on the weather is perceived as indispensable 19, 55. However, we can not expect to have accurate long–term information on future weather conditions. Instead, historical time–series serve as a substitute in capacity expansion planning, and often a single ”representative” sample year is selected for optimisation 41. For example, the International Renewable Energy Agency (IRENA) recommends using $2018$ as a reference year because it represents generation from renewable technologies well on average 31. In the following, we argue that using ”representative” sample years may lead to energy system models (ESMs) that appear sound at first glance, but would fail to meet supply in reality. We then propose a tractable approach to counteract this effect and to make energy systems robust against uncertain conditions during operation, i.e. ensure security of supply.
@@
There is ample evidence in literature that the results of energy system optimisation are sensitive to changes in weather time–series data. Schyska et al. (2021) evaluate the sensitivity of capacity expansion models with regards to multiple sources of uncertainty. They conclude that some years are unsuited as reference years, as using them for optimisation leads to significant misallocation of assets.50 Ruggles et al. (2024) assess how many years of weather data are needed to ensure ESMs are reliable even out of sample, i.e. if a different weather year were to realise. They conclude that between $15$ and $40$ years are required depending on amount of imports available/the acceptable level of loss of load  45. The works by Haddeland et al. (2022) 25, and Staffel et al. (2018)  52, which builds on earlier work by Pfenninger et al. (2017) 39, similarly find the choice of weather years significantly effects renewable generation and thus power output. Not only that, but the effect of the weather increases with increasing share of renewable technologies 8. De Marco et al. (2025) identify energy shortages across Europe and use those to optimize climate–resilient energy systems stochastically 10. Not all research agrees on this; Schlachtenberger et al.(2018) 49 optimise three weather years with hourly data both individually and as one time–series with a resolution of 3h per time step, finding only small variations in total annual cost (TAC) and installed capacities. However, they note that aggregating multiple hours together introduces a smoothing effect that systematically favours PV and underestimate battery and wind generation requirements.
@@
Additionally, there is a fundamental tension between identifying typical years, and build ESMs that are protected against extreme events, i.e. robust. On this, Hilbers et al. (2020) 27 introduce a method of importance subsampling for time–series aggregation to explicitly preserve extreme events in the weather data as an alternative to established ”representative days” clustering approaches. Ryberg’s dissertation 48 and Ryberg et al. (2019) 47investigate the impact of generation lulls in a energy system for a large part of Europe calculating backup capacities required to overcome these. Ruhnau et al. (2022) 46 look into the storage requirements for a renewable–based ESM for Germany using $35$ years of weather data taking consecutive extreme events into account. They conclude that consecutive extreme events increase storage requirements significantly compared to even the most extreme, but singular events. Thus, atypical weather time–series may be particularity well-suited for optimisation because they capture important system behaviours, e.g. dark lulls, that significantly impact ESMs. In this context, Grochowicz et al. (2024) 21 discuss optimising sequential weather years. They use a geometry–based solution approach targeting the solution space. In follow–up work, they use electricity shadow prices to identify difficult weather periods 22. They observe that such difficult weather periods are not just meteorological events, but results of the interplay of meteorology and electricity storage and network structures.
@@
In stochastic optimisation, the notion of using typical, i.e expected, behaviour as a baseline for optimisation is a well known concept. Its usefulness is determined by the value of a stochastic solution, which represents the gap in expected performance of a solution obtained with expected data and one determined by solving a full stochastic optimisation problem  3. In general, this gap may be arbitrarily large. However, research has also shown that we can still extract insights from solutions obtained with with such data  37. Specifically concerning weather robustness, Forghani et al. (2025) propose an intermediate approach where not only a representative but also worst-/best-cost years are used as input data for a stochastic optimisation. Their results show that this reduces loss-of-load to practically zero, at $1\%$ additional cost 14.
An important modelling decision in this context is the value and availability of recourse, i.e. the ability to specify or change parts of the solution if uncertainty realises. In capacity expansion planning, imports constitute such recourse. If arbitrary imports are allowed, any misplanning can be compensated for, if at a cost. We focus on a different setting, the one where imports are limited or energy systems are to be self–sufficient. Here, we use adaptive robust optimisation (ARO) to model limited recourse while still enforcing strict guarantees on security of supply.general solution approaches. To the best of our knowledge, Zeyringer et al. (2018) 58 were the first to assess the effects of weather uncertainty in input data on capacity–expansion planning, looking at an ESM of Grein.They propose using multiple historical time series, optimising over each, and then evaluating the operational costs/supply costs incurred by the proposed energy systems under the assumption that the time–series of a different historical year realises. Then, an ESM with lowest worst–case costs/lowest supply gap amongst the solutions is selected. They find that starting with the ‘’wrong” reference year may lead to misallocation of resources and that modelling with multiple years leads to more consistent results with lower worst–case cost, the latter of which had also been noted before 4. For the US, Dowling et al. (2020) 12 make the case for multi–year modelling to accurately capture long–term storage effects, noting that the cost of variable renewable power systems are especially sensitive to long–duration storage costs. A recent work by Gøtske et al. gøtske2024designingsectorcoupledeuropeanenergy also assesses energy systems based on different weather years. They employ CO2 emitting backup technologies, and analyse structural elements of the respective solutions. Either approach allows to select more suitable ESMs, but it can not assure solutions meet certain supply/demand across all years.
The Two-Stage Robust Setting
@@
To illustrate the setting considered in this work, we begin with an example for an ARO ESM. ARO, alternatively called two-stage robustness, is characterised though a bilevel structure 20 in which a decision maker has to make a set of first stage decisions $x\in\mathcal{X}$, then one of multiple scenarios $u\in\mathcal{U}$ may realise and afterwards the decision maker has the option to react to scenario $u\in\mathcal{U}$ with a scenario-dependent set of decisions $y(u)\in\mathcal{Y}(x,u)$. Equation (ARO) illustrates the structure of a ARO problem. Here, $c$ and $d$ are vectors of first and second stage costs, respectively. The matrix $A$ encodes all information that is not affected by uncertainty, whereas the matrices $B(u)$ and the right-hand side vector $b(u)$ are dependent on the scenario $u$.
@@
For the ESM(s) considered in this work, the first stage decisions correspond to the capacity expansion planning, e.g. investment in solar power plants, power lines, storage units, and the second stage corresponds to operational decisions, e.g. how much energy is to be stored in which battery at what time. If energy imports are allowed, these will also be encoded in $\mathcal{Y}$. The scenario $u$ corresponds to a specific, continuous time period. Table 1 illustrates this setting. Here, we consider different locations $A,B,C$ for a solar power plant, whose production is subject to changing weather patterns during the time periods $u_{1},u_{2}$.
@@
Table 1: Investment problem: Chose number of solar power plants to build on locations $A,B,C$. Values indicate expected energy supply in $\frac{\mathrm{GWh}}{\mathrm{time~period}}$ per investment.
@@
The optimal solution depends on what type of recourse is available and how much supply is needed. For example, if we assume similar cost at all locations and no/expensive imports, investing in location $B$ is the most cost effective for meeting a demand of $75\frac{\mathrm{GWh}}{\mathrm{time~period}}$. However, if the total demand were $200\frac{\mathrm{GWh}}{\mathrm{time~period}}$, it is more effective to invest in $A$ and $C$ instead of $B$.
@@
Returning to our example in Table 1, the approach by Zeyringer et al. would not be able to find the optimal robust solution for a total demand of 7575MWh/year, which is to invest in location $B$. This is because if the future is known, i.e. we assume either $u_{1}$ or $u_{2}$ to realise, the optimal investment is always to either invest in location $A$ or location $B$. An alternative approach was put forward by Caglayan et al. (2019) 5, who noted variations of up to $20\%$ variations in TAC, and significant differences in energy system designs when optimising an integrated electrical/hydrogen ESM for Western Europe with different weather years. They propose an iterative approach, where they begin with optimising individual years, but then average their designs and take the average as a lower bound for installed capacities in a new round of optimisation 6. Since the (average) capacities installed during sequential algorithm iterations will be non-decreasing and bounded, convergence and thus a feasible solution will be guaranteed in most cases. In our example, the algorithm would enforce capacity >0>0 for location A and C, which would eventually lead to a feasible but unnecessarily expensive solution that uses both A and C.
A Novel Algorithm for Computing Robust ESMs
@@
The proposed algorithm consists of four main steps, as shown in Figure 1. As its input, the algorithm takes a set of $n$ times series, one of which is designated as a starting/reference year. We begin with solving the capacity expansion planning problem for this year (CAPEX), then we evaluate the performance of the proposed investment in all $n$ years by solving the corresponding unit commitment problem (UC), which formalises the validation proposed in previous works  58, gøtske2024designingsectorcoupledeuropeanenergy.
To ensure feasibility, existing approaches allow for energy imports. We consider the robust setting where imports are bounded/not available, i.e. the ESM needs to be self-sufficient. Based on the previous step, we identify supply gaps, which we denote by $\delta$. The underlying ESM for Germany for $2045$ utilizes renewable technologies only and includes PV, wind, Li–ion batteries, electrolyzer, hydrogen salt caverns, hydrogen combined cycle gas turbines as well as electricity transmission and hydrogen pipelines. The models are described in detail in the methods section for both gurobipy and the ETHOS.FINE. If the supply gaps are sufficiently small, we terminate. Otherwise, we use the information encoded in $\delta$ to iteratively make solutions more robust. For that, modifications (MOD) are applied to the optimisation problem by either adding artificial demand MOD 1 / Demand increase, substituting parts of the time series data MOD 2 / Synthetic time–series or iteratively adding both demands and lazy valid inequalities MOD 3 / Combine, before reoptimising CAPEX to evaluate the loss of load. Detailed definition of the modifications can be found in the methods section.
@@
Start: Solve CAPEX for reference year Solve UC for all $n$ years Small loss of load? Apply MOD1-3 to CAPEX Reoptimise modified CAPEX End: Robust energy system design yesno
@@
The main contributions of our work include not only insights derived for the specific ESM, and an algorithm for solving CAPEX problems to find solutions that are robust to variations in operational conditions across the $n$ years, but foremost a systematic comparison of different approaches for designing robust ESMs. Based on this, we recommend three working modifications that all lead to robust energy systems systems.
@@
In the following, we first compare solutions derived from individual years’ data and assess their (lack of) robustness. Then, we modify them to become more robust, and evaluate the features of robust solutions. Note that we provide results for two different models, since some modifications require custom callbacks/lazy constraints which are hard to integrate into modelling frameworks. Thus, we use one model directly coded in gurobipy that allows for all modifications but is simplified, see Appendix A, and a more realistic model implemented in the ETHOS.FINE framework.
Note that we also only use a single demand time-series from $2012$ for comparison. This significantly simplifies evaluation of the algorithms as otherwise, loss of load and installed would need to be considered relative to each year’s demands. However, in most practical settings where demands and weather are linked, e.g. when employing significant electric or hydrogen-based heating, the paired demand time-series ought to be used.
Optimal Capacities Strongly Depend on Yearly Weather Data
The TAC for energy systems for the $38$ node model of Germany within ETHOS.FINE for the $40$ different years of time–series data deviates around an average of 106106bn€ with cost between 96.496.4bn€  and 113.6113.6bn€ annually, which equals $9\%$ to $7\%$ compared to the default reference year $2018$ recommended by IRENA 31. The results of optimising each year independently are given by Figure 2. While the variations in overall TAC are limited, the energy system designs show substantial deviations. The cost shares of single technologies across the $40$ different single years vary by $69\%$ for hydrogen pipelines, $57\%$ for hydrogen salt caverns, $53\%$ for CCGT, $44\%$ for Li–ion batteries, $40\%$ for PV, $38\%$ for electrolysers, $22\%$ for the electricity grid, and $20\%$ for onshore wind, making it challenging to draw recommendations for planned capacity expansion for future energy systems.
Figure 2: TAC comparison by technology for ESMs optimised from 1980–2019 aggregated for a $38$ node Germany model set up in ETHOS.FINE.
@@
Figure 3: TAC comparison by technology for ESMs optimised from 198020191980-2019 on a single node Germany model set up using gurobipy.
@@
In summary, even small input data differences may lead to large variations in investment into different technologies. For example, offshore wind is not utilised at all in some years, but makes up over $13\%$ of the TAC in $2014$. Note that the single node model is not directly comparable to the ETHOS.FINE model, as it contains various simplifications. As such, solutions are significantly cheaper averaging 55.455.4bn€ with between $52.1$ and 59.759.7bn€ annually, i.e. $6\%$ to $7\%$.
Systems Optimised on Single Years Cause Supply Gaps
Feasibility testing shows that all $40$ energy system designs solely based on one year’s time–series lead to supply gaps in multiple other years. This means that none of the system designs are robust. However, we are not only interested in whether there is any loss of load, but also in the magnitude of said loss of load. Figure 4 shows the amount of load shedding when testing the feasibility of one of five reference years used in ETHOS.FINE during the other $39$ years.
Figure 4: Load shedding when testing the feasibility of the $5$ selected reference years in ETHOS.FINE. Only the years and the months from October to March are included, where load shedding occurs. For comparison, hourly load in the demand data we use is roughly 100100GW.
Two trends can be observed: On the one hand, when the feasibility is tested for $1987$ and $2010$, load shedding occurs mostly in the first half of February. This is caused by demand spikes in early February due to the low temperatures in that time of the year, which the electricity demand was based on. At the same time, the energy systems optimised for $1987$ and $2010$ are characterised by the highest amounts of PV and wind capacities installed while having insufficient backup capacities.
On the other hand, the energy system of the reference years $2009$ has increased backup capacities, therefore suffering less during cold darklull periods, but does not have enough PV and wind capacities to produce enough hydrogen and therefore has load shedding more evenly distributed over the months from October to March. The other two reference years $1996$ and $2013$ are in the middle of these two case.
@@
Figure 5 illustrates the reasons for that by showing three uncritical (top row) periods with sufficient supply as well as three critical (bottom row) time periods with significant supply gaps. During critical time periods electricity generation from PV and Wind is low, providing less than 50% of electricity demand. The remaining electricity demand needs to be fulfilled by backup capacities, i.e. $H_{2}$ CCGT. The periods have varying duration and they are identified via clustering and feasibility testing. Uncritical time periods are characterised by high availability of PV or onshore wind or both, while critical ones are characterised by low availability of PV and low to negligible onshore wind combined with high demand. Offshore wind plays only a minor role due to its limited utilisation. The most critical time period takes place in $1994$, see Figure 5f). Here, wind and solar supply indicate a dark lull. Combined with the cold period identified in the electricity data, this constitutes a cold dark lull.
@@
Figure 5: Six time periods from the $40$ weather years and 11 year of future electricity demand data for Germany. The electricity demand is normalised to prevent overweighing; the weather data is aggregated. The upper three diagrams represent non–critical time periods, the lower three critical ones. Subfigure a) is a typical summer period with high PV availability and low electricity demand due to low heating requirements. In b) and c), typical autumn and winter period are shown. They are characterised by low availability of PV, but ample wind power to supply sufficient electricity. Note the increased electricity demand due to increased heating required. In d), a typical dark lull is characterised by low availability of PV and negligible amounts of wind, which coincides with high electricity demand due to increased heating. Subfigure e) shows an elongated dark lull period. Low availability of both PV and wind combined with increased electricity demand lead to overall difficult period requiring large amounts of hydrogen to be burned in the energy system. The last graphic f) shows the most critical period in the 40 years of data. Negligible wind combined with low availability of PV coincide with the highest electricity demand in the data due to high heating demand during an extreme cold spell in all of Germany.
For the gurobipy implementation, we get similar results: The minimal annual supply gap across all years is 1.1701.170GWh for $1987$, which is about $0.13\%$ of the total annual electricity demand. Additionally, during shorter time-periods the supply gap can reach up to half of the required electricity demand. In comparison, optimising using $1990$’s weather data leads to a peak supply gap of 69.50069.500GWh, i.e. $8\%$, if the weather of $1987$ were to realise.
In summary, this strongly supports the results by Cagalayan et al. 5, 6 – weather patterns matter and ignoring in energy systems with significant renewables may lead to systems that have large supply gaps under anything but the most optimal conditions, which we cannot simply assume to be covered through imports.
Effect of Modifications on Installed Capacities
@@
All modifications lead to increased investment in CCGT, salt caverns and electrolysers, although the latter is less pronounced for MOD 1 / Demand increase, as seen in Table 3. This is expected, as CCGT can provide electricity when the generation is low and demand is high and installing more CCGT also requires more H2 infrastructure such as salt caverns for storage and electrolysers for H2 conversion.